# Quantum gravity cannot be both consistent and complete

### v1

@@ V Application to Physical Problems
We now discuss some applications of these results, starting from the black hole information paradox Almheiri2021 . It could be that the information stored a black hole associated with Planck scale physics, where quantum fluctuations and traditional notions of geometry appear to break down, cannot be microscopically algorithmically computed due Chaitin’s incompleteness theorem, as alluded to above. In this case, the emergence of smooth spacetime from Planck-scale physics would occur through thermalization, where microscopic quantum gravitational degrees of freedom collectively evolve into a macroscopic geometric state. This could be beyond algorithmic computation, as in general it has been explicitly proven that determining whether a given system thermalizes is computationally undecidable Shiraishi2021 .
@@
Also, using an explicit construction of quantum spin models whose phase diagrams encode uncomputable problems, it has been demonstrated that no general algorithm can fully determine these diagrams bausch2021uncomputability , implying intrinsic computational limits to predicting phase transitions and understanding complex quantum matter. Now as the quantum spin model and LQG are closely related mathematically Feller:2015 , this may have important implications for LQG. So, it is possible that the the full phase diagram of LQG, like quantum many-body systems, is fundamentally not computable.
These results suggest that certain truths lie beyond the computational domain of any consistent and sufficiently expressive theory of quantum gravity. Note that the principle of sufficient reason states that every fact has a reason, which forms the basis of science itself amijee2021principle ; leibniz1996discourse . The absence of an algorithmic explanation does not contradict this principle; rather, it suggests that while some explanations may be expressed in algorithmic terms, others reside in a non-algorithmic understanding, which still provides a sufficient explanation for the phenomena in question.
@@
Gravitationally induced objective collapse, which addresses the measurement problem in quantum mechanics using an objective physical process tied to gravitational effects, also exemplifies the application of non-algorithmic understanding Penrose1996 ; Diosi1987 . In this framework, measurement occurs independently of an observer, and is neither random nor computational. In this regard, it has been explicitly demonstrated that quantum logic is inherently undecidable due to incompleteness theorems vandenNest2008measurement ; lloyd1993quantum . So, an approach based on non-algorithmic understanding Penrose1996 ; Diosi1987 produces a better physical explanation of the measurement problem. It is well established that there are limitations to computations in various problems of physics peraleseceiza2024undecidabilityphysicsreview . Here we have argued that these similarly imply it impossible for quantum gravity based on computations to be both consistent and complete.
To address the inherent limitations of formal systems in quantum gravity, a meta-theory of everything $\mathcal{M}_{ToE}$ is needed, which contains non-algorithmic understanding. This meta-theory of everything would extend the computational-based formal system $\mathcal{F}_{QG}$ by introducing a truth predicate $T(x)$, enabling the recognition of truths beyond formal provability. Specifically, $\mathcal{M}_{ToE}$ would allow for the acceptance of Gödel statements $G$ as true, the resolution of undefinable truths, and the acknowledgment of high-complexity statements $S$ with $K(S)>K_{\mathcal{F}_{QG}}$. This meta-theory would satisfy: $T(G)=1\quad\text{for }G\text{ true but unprovable in }\mathcal{F}_{QG}.$ By transcending the algorithmic limits of $\mathcal{F}_{QG}$, $\mathcal{M}_{ToE}$ would provide a framework to explore truths that lie beyond computation. Inspired by principles such as Novikov self-consistency, and similar principles based on non-algorithmic understanding, this approach could integrate rigorous computations with non-algorithmic insights. It could potentially resolve challenges like the black hole information paradox and the complexities of Planck-scale physics.
Considerations of quantum gravity have suggested that physics should evolve from a focus on ‘it’ (i.e. matter fields moving in spacetime) to a focus on ‘bit’, i.e., information being more fundamental. The considerations we present here show that neither ‘it’, nor ‘bit’ is fundamental. Namely, a non-algorithmic understanding must operate at a fundamental level in any theory of everything.

### v3

@@ V Application to Physical Problems
We will now discuss some potential applications of such a structure beyond computation in quantum gravity. It should be noted that we will not explicitly prove the undecidability of these structures, but will instead provide physical arguments as to why they are expected to be undecidable. We will also mention physical systems in which specific properties have been shown to correspond to mathematical undecidable statements peraleseceiza2024undecidabilityphysicsreview . We notice however that, in most cases, these results have been obtained via the halting problem of a universal Turing machine or through other mathematical systems embedding or simulating a universal Turing machine. It is possible to show that the undecidability of the halting predicate and Gödel’s incompleteness phenomena are extensionally equivalent. By encoding a Gödel-style self-referential sentence as the source code of a program, one constructs an arithmetic statement whose provability within any consistent, computably enumerable theory would decide whether that program halts. If the theory could resolve every such statement, it would effectively compute the universal halting problem, contradicting Turing’s diagonal argument. Conversely, the inability of the theory to settle the program’s behavior mirrors its inability to prove its own consistency. Hence Gödel’s incompleteness theorems follow almost immediately from the halting problem, affirming that no sufficiently expressive formal system can simultaneously prove its consistency and decide all halting instances Calude2021Incompleteness . As previously noted, undecidability arises naturally within Selective Actualism. In the following, however, we employ physical reasoning to argue that it also holds within the framework of Universal Actualism, and that computation alone is insufficient to fully account for actual physical phenomena.
We start our analysis from the black hole information paradox Almheiri2021 . It could be that the information stored a black hole associated with Planck scale physics, where quantum fluctuations and traditional notions of geometry appear to break down, cannot be microscopically algorithmically computed due Chaitin’s incompleteness theorem, as alluded to above. In this case, the emergence of smooth spacetime from Planck-scale physics would occur through thermalization, where microscopic quantum gravitational degrees of freedom collectively evolve into a macroscopic geometric state. This could be beyond algorithmic computation, as in general it has been explicitly proven that determining whether a given system thermalizes is computationally undecidable Shiraishi2021 .
@@
Also, using an explicit construction of quantum spin models whose phase diagrams encode uncomputable problems, it has been shown that no general algorithm can fully determine these diagrams bausch2021uncomputability , implying intrinsic computational limits to predicting phase transitions and understanding complex quantum matter. Now as the quantum spin model and LQG are closely related mathematically Feller:2015 , this may have important implications for LQG. So, it is possible that the the full phase diagram of LQG, like quantum many-body systems, is fundamentally not computable.
These results suggest that certain truths lie beyond the computational domain of any consistent and sufficiently expressive theory of quantum gravity. Note that the principle of sufficient reason states that every physical fact has a reason, which forms the basis of science itself amijee2021principle ; leibniz1996discourse . The absence of an algorithmic explanation does not contradict this principle; rather, it suggests that while some explanations may be expressed in algorithmic terms, others reside in a non-algorithmic understanding, which still provides a sufficient explanation for the phenomena in question.
@@
Gravitationally induced objective collapse, which addresses the measurement problem in quantum mechanics using an objective physical process tied to gravitational effects, also exemplifies the application of non-algorithmic understanding Penrose1996 ; Diosi1987 . In this framework, measurement occurs independently of an observer, and is neither random nor computational. In fact, it has been explicitly demonstrated that even quantum logic is inherently undecidable due to incompleteness theorems Fritz2021 ; Eisert2012 ; vandenNest2008measurement ; lloyd1993quantum . So, an approach based on non-algorithmic understanding Penrose1996 ; Diosi1987 produces a better physical explanation of the measurement problem. It is well established that there are limitations to computations in various problems of physics peraleseceiza2024undecidabilityphysicsreview . Here we have argued that these similarly imply it impossible for quantum gravity based on computations to be both consistent and complete. Here, it is important to note that even if some specific phenomena discussed here are not undecidable, the general arguments presented imply that there will definitely be some undecidable phenomena. Hence, this problem will inevitably arise, and we need to go beyond a computational framework to address it.
To address the inherent limitations of formal systems in quantum gravity, a meta-theory of everything $\mathcal{M}_{ToE}$ is needed, which contains non-algorithmic understanding. This meta-theory of everything would extend the computational-based formal system $\mathcal{F}_{QG}$ by introducing a truth predicate $T(x)$, enabling the recognition of truths beyond formal provability. Specifically, $\mathcal{M}_{ToE}$ would allow for the acceptance of Gödel statements $G$ as true, the resolution of undefinable truths, and the acknowledgment of high-complexity mathematical statements $S$ with $K(S)>K_{\mathcal{F}_{QG}}$. This meta-theory would satisfy: $T(G)=1\quad\text{for }G\text{ true but unprovable in }\mathcal{F}_{QG}.$ By transcending the algorithmic limits of $\mathcal{F}_{QG}$, $\mathcal{M}_{ToE}$ would provide a framework to explore truths that lie beyond computation. Inspired by principles such as Novikov self-consistency, and similar principles based on non-algorithmic understanding, this approach could integrate rigorous computations with non-algorithmic insights. It could potentially resolve challenges like the black hole information paradox and the complexities of Planck-scale physics.
Considerations of quantum gravity have suggested that physics should evolve from a focus on ‘it’ (i.e. matter fields moving in spacetime) to a focus on ‘bit’, i.e., information being more fundamental. The considerations we present here suggest that neither ‘it’, nor ‘bit’ is fundamental. Namely, a non-algorithmic understanding should operate at a fundamental level in any theory of everything.